# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v1

@@ 2 Related Work
Table 2: Contradiction types from (de Marneffe et al., 2008). Contradiction covers a broader scope than negation.
Negations in generative language models.
Recent research on negation understanding has primarily focused on bidirectional models like BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) due to their strong performance on NLU and negation detection tasks. However, with the rise of generative foundation models such as GPT (Radford et al., ) and LLaMA (Touvron et al., 2023), attention has shifted toward evaluating their negation handling. Studies have found that these models often exhibit positive bias and struggle with producing or interpreting negated statements (Truong et al., 2023; Chen et al., 2023; García-Ferrero et al., 2023). While benchmarks, such as CONDAQA (Ravichander et al., 2022) and ScoNe (She et al., 2023), expose these limitations, robust evaluation resources tailored to generative models remain scarce.
Building upon these prior studies, this paper evaluates whether generative models can understand negation in complex sentences and distinguish subtle semantic differences beyond surface-level patterns.
3 Definition and Scope of Negation
Although negation has been widely studied in NLP, its definition and scope remain loosely specified, with most studies focusing on identifying negation cues or confusing negation with contradiction. In this work, we aim to refine the definition of negation by examining its semantic boundaries, distinguishing it from related but distinct phenomena such as contradiction, and characterizing the types of meaning reversal.
3.1 Typology of Negation
Negation is a core semantic and syntactic operation in natural languages expressing a proposition’s denial, contradiction, or absence. In formal logic, it reverses the truth value of a proposition: if $P$ is true, then $\neg P$ (which means the negation of $P$) is false, and vice versa. Semantically, negation introduces oppositions, often to a corresponding affirmative proposition (Horn and Wansing, 2025).
Negation can be categorized along several dimensions, including scope, form, and target (see Table 1). By scope, it may affect the entire clause (clausal negation) or just a part of it (subclausal negation). In terms of form, it can appear as bound morphemes, such as prefixes and suffixes (morphological negation), or as separate syntactic elements like "not" or "never" (syntactic negation). Finally, depending on its target, negation can apply to the verb (verbal negation) or to other elements in the sentence (non-verbal negation)  (Zanuttini, 2001; Miestamo, 2007; Truong et al., 2022; Kletz et al., 2023).
3.2 Negation and Contradiction
Negation and contradiction, closely related concepts, are often conflated in NLP research (Jiang et al., 2021). Contradiction refers to the incompatibility of two propositions: they cannot be both true simultaneously. Although negation frequently serves as a primary mechanism to create contradictions by reversing the truth value of a proposition, contradictions may also arise through antonymy, numeric mismatch, structural and lexical differences (more details are in Table 2(de Marneffe et al., 2008). Previous studies have often overlooked the possibility of contradictions existing independently of explicit negation. Recognizing this gap, we specifically examine the ability of LLMs to differentiate between negations and non-negated contradictions, highlighting the nuanced semantic distinctions involved.
3.3 Standard Negation
In this paper, we specifically examine standard negation: the prototypical form of negation applied to the main declarative verbal clause. Standard negation involves negating the main verb in a main clause, where the main verb expresses the main action of a clause, and the main clause itself can independently form a complete sentence. This definition excludes negation found in subordinate clauses, which are clauses dependent on a main clause  (Miestamo, 2000). This paper specifies the standard negation as reversing the truth value of the main predicate (verbal phrase). Formally, if a main predicate is denoted as $P$, standard negation corresponds precisely to $\neg P$ (i.e., the logical negation of $P$).
Standard negation can encompass various dimensions, as described in Table 1. Specifically, it includes clausal negation, which affects the entire sentence, and verbal negation, which explicitly targets the main verb. In terms of form, standard negation can be realized either syntactically or morphologically.
Syntactic negation typically involves inserting negation particles (e.g., "not") to directly negate the main predicate. On the other hand, morphological negation is more limited, applying only when the antonym of the main predicate fully encompasses its mutually exclusive semantic space (e.g., "be alive" vs. "be dead"). Thus, morphological negation qualifies as standard negation only in cases involving complementary antonyms, which represent absolute binary oppositions (e.g., "true" vs. "false" and "possible" vs. "impossible").
In contrast, other types of antonyms, such as gradable antonyms—words that express opposite meanings along a spectrum of quality (e.g., "happy" vs. "unhappy/sad/depressed")—and relational antonyms—words expressing opposite relational roles (e.g., "buy" vs. "sell")—do not strictly reverse truth values (Lehrer and Lehrer, 1982). These examples fall under contradiction rather than the standard negation discussed in this paper.
The above definitions and scope of negation provide the foundation for constructing our benchmark dataset, which we describe in detail in the following section.
4 Thunder-NUBench Dataset
Figure 1: Dataset generation process.
We construct the Thunder-NUBench dataset based on two datasets: (1) HoVer dataset, which is designed for multi-hop fact extraction and claim verification based on Wikipedia articles (Jiang et al., 2020), (2) Wikipedia Summary dataset, which contains concise article titles and summaries extracted from English Wikipedia (Scheepers, 2017). We select these datasets as our base corpora because their factual content and complex sentence structures make them well-suited for creating a dataset to understand standard negation in long sentences.
4.1 Dataset Generation
The overall process for dataset generation proposed in this paper is illustrated in Figure 1. After extracting sentences from the two sources and preprocess them, we construct two types of datasets: (1) a sentence-negation pair dataset, which includes only standard negation examples generated, and (2) a multiple choice dataset, which covers four categories, systematically constructed through a combination of manual authoring and automated generation. All data is then verified and refined through human review, with a strict protocol to ensure that no author reviews data they generated to ensure quality and consistency. Further details of each step are described below.
Sentence-negation pair dataset.
We begin by randomly sampling sentences labeled as "supported facts" from the HoVer dataset. Since the original data often contains grammatical errors, we utilize OpenAI’s API (OpenAI, 2025) to automatically correct these issues. In cases where the selected text consists of multiple sentences, we merge or split them as needed to ensure that each example is a single sentence, aligning with our sentence-level task objective (see Appendix D for details). Each sentence is manually negated according to the standard negation criteria described earlier, followed by a thorough review process.
Multiple choice dataset.
To construct the multiple-choice dataset, we first segment the "summary" column of the Wikipedia Summary dataset into individual sentences, which often contain multiple sentences in a single entry. To focus on the challenges of negation in complex sentences, we filter out sentences that are too short. We generate multiple candidate options for each selected sentence, including standard negation, local negation, contradiction, and paraphrase.
Negation examples are manually written because LLMs often fail to generate accurate negations, frequently producing subclausal negations when standard negation is required, or generating incorrect local negations even when explicitly prompted. As a result, automated generation is not used for these cases. In contrast, non-negated contradiction and paraphrase examples are first generated automatically using carefully designed prompts with the OpenAI API. All data are further reviewed and refined by the authors.
Type Structure Explanation Local Negation Example
Relative Clause
Negation
A relative clause is a type of dependent clause that gives
extra details about a noun or noun phrase in the main sentence.
It usually begins with a relative pronoun such as
who, which, that, whom, or whose.
The man who owns the car
is my neighbor. $\rightarrow$ The man
who does not own the car is
my neighbor.
Participle Clause
Negation
A participle clause is a type of dependent clause that begins with
a participle (a verb form ending in -ing or a past participle).
It acts like an adverb, giving extra details about the main
clause, often showing time, reason, result, or sequence of actions.
Walking through the park, she
found a lost wallet. $\rightarrow$ Not
walking through the park, she
found a lost wallet.
Adverbial Clause
Negation
An adverbial clause is a dependent clause that acts like an adverb,
modifying a verb, adjective, or adverb. It gives information such
by subordinating conjunctions like because, although, or while.
She stayed inside because it was
raining. $\rightarrow$ She stayed inside
because it was not raining.
Compound
Sentence
with
Partial Negation
A compound sentence consists of two or more main clauses
joined by coordinating conjunctions such as and, but, or or.
If only one of these clauses is negated, the negation applies
only locally to that clause.
He submitted the report and
attended the meeting. $\rightarrow$ He
submitted the report and did
not attend the meeting.
Table 3: Typology of Local Negation.
4.2 Human Curation Rules
Below, we describe the principles for manual generation and dataset review. Additional details and examples are provided in Appendix E.
Standard negation.
The rules for standard negation are as follows:
  • The standard negation is intended to reverse the sentence’s overall meaning and is achieved by negating the main clause’s main verb, reversing the sentence’s truth value. All other elements of the main clause (subject, object, temporal context, etc.) are preserved unchanged.
  • Standard negation is implemented by inserting negative particles such as "not" into the main verb or substituting the main verb with its complementary antonym, when appropriate.
  • The main verb may be replaced with a synonymous verb only if the overall meaning remains strictly identical. Other components may be paraphrased with synonyms as long as the tense, sentence structure, and meaning remain equivalent. Standard negation is then applied to the paraphrased sentence.
  • For compound sentences joined by coordinating conjunctions (and, or, and but), negation follows logical rules, such as De Morgan’s laws (e.g., A and B $\rightarrow$ $\neg$¬A or $\neg$¬B, where A and B are clauses). For unnatural outputs, sentences may be split for fluency, as long as logical negation is preserved.
Local negation.
We define local negation as a negation that targets a verb phrase outside the main clause. This work applies local negation to four types of sentence structure: relative clause, participle clause, adverbial clause, and compound sentence (see Table 3 for more details). In compound sentences, standard negation requires that all main clauses be negated to achieve sentence-level negation; if only a subset of clauses is negated, it is considered local negation. The mechanism for constructing local negation follows that of standard negation, but the negation is applied only to the intended subpart rather than the entire main clause.
This design allows us to test whether models can distinguish between full-sentence negation and local negation. Although the scope of local negation is restricted, it still contains explicit negation cues (e.g., "not"), which may mislead models that rely on shallow cue detection rather than deeper semantic understanding.
Contradiction.
Contradiction examples in this work are constructed using mechanisms such as antonymy, numeric changes, or structural alterations, as long as the resulting sentence cannot be true at the same time as the original and does not simply apply standard or local negation. Unlike standard negation, which reverses the truth value of the main predicate, contradiction can arise from modifying adjectives, quantities, named entities, or other semantic elements. Both negation and contradiction can involve antonyms; however, only complementary antonyms are permitted for standard and local negation, while gradable or relational antonyms are allowed for contradiction. During validation and review, authors ensure that no pair of original and contradictory sentences can be simultaneously true.
It is important to note that standard negation is a subset of contradiction: every negation is a contradiction, but not every contradiction is a standard negation. The goal of this category is to assess whether models can reliably distinguish standard negation from other forms of contradiction, as both alter the meaning of a sentence, but standard negation reverses the entire proposition, whereas contradiction, as defined in this work, does not necessarily do so.
Paraphrase.
A paraphrase rewrites the sentence using different wording or structure while preserving the original meaning. No additional information may be introduced, and the main verbs and core content must remain unchanged. Identical or near-identical sentences of the original sentence, which often occur in automatically generated paraphrases, are carefully screened and omitted as well.
The reason for including paraphrase examples is to test whether models incorrectly interpret sentences with different surface forms (e.g., synonyms, rephrased structures) as having reversed meanings. This allows us to examine the robustness of language models in distinguishing genuine reversals from similar meaning, a distinction that has been highlighted as a challenge in previous research on distributional semantics (see Section 2).
Dataset Split Count
Sentence-Negation Pair Dataset Train 3,772
Multiple Choice Dataset Validation 100
Test 1,002
Total 4,874
Table 4: Thunder-NUBench Dataset Statistics.
4.3 Dataset Statistics
The final dataset consists of a sentence-negation pair training set and a multiple-choice evaluation set (see Table 4). The multiple choice dataset presents each original sentence with four options: standard negation (choice1, always the answer), local negation (choice2), contradiction (choice3), and paraphrase (choice4).
To construct the validation set of the multiple-choice dataset, we first select 100 examples whose Wikipedia page indices are unique within the dataset, preventing any duplication with the test set. We also matched the distribution of local negation types to the overall dataset as closely as possible, ensuring that the validation set serves as a representative subset. Thunder-NUBench is available online. [^1] https://huggingface.co/datasets/thunder-research-group/SNU_Thunder-NUBench
5 Experiments
We conduct experiments on Thunder-NUBench using an instruction-based prompt that explicitly includes logical rules (as illustrated in Listing 1; more details of prompt selection in Appendix J).
1"Logically negate the sentence below. If the sentence includes A and B’, use not A or not B’. If it includes A or B’, use not A and not B’. Also apply not or use complementary antonyms on the main verb(s) of the entire sentence.
2Sentence: {doc[’sentence’]}
3Negation:"
Listing 1: Instruction-based prompt format.
Zero-shot and few-shot.
For each model, we evaluate both zero-shot and few-shot settings using the Language Model Evaluation Harness (Gao et al., 2024). In the few-shot scenario, we use examples from the validation set as in-context demonstrations. In few-shot experiments, results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and five examples of validation set data (5 shots). We report performance on the test set for each model and prompt configuration. All the results are provided in Appendix M.
Baseline
SFT on
Thunder-NUBench
zeroshot
fewshot
(±SD)
zeroshot
Qwen2.5-3B acc 0.495
0.600
(±0.009)
0.747
acc_norm 0.526
0.622
(±0.009)
0.753
Qwen2.5-3B- Instruct acc 0.667
0.758
(±0.006)
0.765
acc_norm 0.686
0.771
(±0.006)
0.775
Qwen2.5-7B acc 0.499
0.623
(±0.009)
0.820
acc_norm 0.521
0.633
(±0.012)
0.825
Qwen2.5-7B- Instruct acc 0.545
0.700
(±0.004)
0.783
acc_norm 0.576
0.713
(±0.002)
0.784
Average acc 0.551 0.670 0.779
acc_norm 0.577 0.685 0.784
Table 5: Evaluation results of Qwen2.5 models across different settings: baseline zero-shot, baseline few-shot (5-shot), and zero-shot after supervised fine-tuning (SFT) on Thunder-NUBench. Both accuracy (acc) and normalized accuracy (acc_norm) are reported. Few-shot results are averaged over five random seeds with the standard deviation in parentheses.
Error rate (1-acc) Incorrect Choice Distribution choice2 confusion rate
choice2
choice3
choice4
relative_part
confuse
pp_part
confuse
compound_part
confuse
adverb_part
confuse
Llama-3.1-8B baseline zeroshot 0.523 70.80 22.90 6.30 29.35 34.83 53.43 34.29
fewshot 0.357 81.01 16.20 2.79 22.53 21.38 47.29 29.52
After SFT zeroshot 0.149 80.54 16.11 3.36 9.56 10.34 18.77 9.52
gemma-7b baseline zeroshot 0.508 68.17 28.88 2.95 28.67 33.10 46.93 35.24
fewshot 0.382 83.81 14.62 1.57 26.28 23.45 49.10 38.10
After SFT zeroshot 0.153 77.78 20.26 1.96 9.56 10.34 19.49 6.67
Qwen2.5-7B baseline zeroshot 0.501 66.14 30.08 3.78 31.40 33.45 40.07 30.48
fewshot 0.387 74.23 24.23 1.55 24.91 24.48 41.52 27.62
After SFT zeroshot 0.180 85.56 13.33 1.11 14.33 15.52 20.22 10.48
Mistral-7B-v0.3 baseline zeroshot 0.521 72.03 22.61 5.36 26.96 34.83 58.12 33.33
fewshot 0.372 79.09 18.23 2.68 22.87 22.76 48.01 27.62
After SFT zeroshot 0.161 83.85 13.66 2.48 8.87 11.03 25.27 6.67
Table 6: Incorrect choice distribution and confusion analysis in negation benchmark across various 7-8B size pretrained models. Few-shot results are reported with a fixed seed (1234) to keep error patterns clear because averaging over multiple seeds could make them harder to interpret.
SFT.
We perform Supervised Fine-Tuning (SFT) using the LLaMA-Factory framework (Zheng et al., 2024) on the Sentence-Negation Pair Dataset from Thunder-NUBench. The data is formatted in the Alpaca instruction style (Taori et al., 2023). For parameter-efficient training, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank of 8, targeting all linear layers. Fine-tuning is conducted for three epochs with a batch size of 1, a gradient accumulation step of 8, cosine learning rate scheduling, and bfloat16 precision. After SFT, we evaluate zero-shot performance to directly measure the model’s ability to generalize from instruction tuning without the influence of in-context examples. All the results are provided in Appendix N.
Benchmarking with Thunder-NUBench.
Table 5 presents evaluation results for Qwen2.5-3B, Qwen2.5-3B-Instruct, Qwen2.5-7B, and Qwen2.5-7B-Instruct models (Qwen et al., 2025) in three settings: zero-shot baseline, few-shot baseline, and zero-shot after SFT using Thunder-NUBench.
Instruction-tuned models consistently outperform their pretrained counterparts. Few-shot prompting significantly improves performance, highlighting the benefit of concrete examples. Supervised fine-tuning on Thunder-NUBench further boosts accuracy, with pretrained models showing the most significant gains. Notably, Qwen2.5-3B-Instruct performs exceptionally well in zero-shot and few-shot settings, even outperforms larger models, suggesting strong alignment with the logical reasoning demands of Thunder-NUBench.
Negation understanding analysis.
We analyze model errors to assess their ability to distinguish standard negation from similar semantic variants. Each local negation subtype in our dataset is explicitly labeled according to its sentence structure: relative clauses (relative_part), participle clauses (pp_part), compound sentences (compound_part), and adverbial clauses (adverb_part). To identify which subtypes are most often confused with standard negation, we compute the confusion rate, defined as the proportion of examples within each subtype where the model incorrectly selects the local negation option (choice2) instead of the correct standard negation (choice1).
For instance, among 1,002 test examples, 290 of choice2 are labeled as pp_part; if the model erroneously chooses choice2 in 29 of these, the confusion rate for pp_part is 10%. For all the details, please see Appendix P.
We conduct a comparative analysis across four 7-8B scale pretrained language models (LLaMA-3.1-8B, Gemma-7B, Qwen2.5-7B, and Mistral-7B-v0.3) (Grattafiori et al., 2024; Team et al., 2025; Qwen et al., 2025; Jiang et al., 2023) under three evaluation settings: baseline (zero-shot and few-shot) and zero-shot after SFT. All models show a consistent tendency to incorrectly select local negation (choice2), even though SFT reduces overall errors, showing that distinguishing local from full-sentence negation remains difficult. Confusion rate is especially high for compound sentence structures, highlighting specific areas where models systematically struggle with negation understanding.
6 Conclusion
In this work, we introduce Thunder-NUBench, a benchmark designed to evaluate large language models(LLMs)’ sentence-level understanding of negation. Unlike prior sources that isolate negation in simple syntactic contexts, Thunder-NUBench incorporates structurally diverse alternatives, such as local negation, contradictions, and paraphrases, to assess models’ deeper semantic comprehension. Our evaluations demonstrate that while LLMs benefit from few-shot prompting and fine-tuning, they struggle to distinguish standard negation from closely related semantic variations. Thunder-NUBench provides a valuable diagnostic tool for analyzing models’ limitations regarding negation understanding, and serves as a robust benchmark for future research.
Limitations
Ethical Considerations
The Thunder-NUBench dataset is released under the CC BY-NC-SA 4.0 license, ensuring transparency, reproducibility, and accessibility for future research. We believe our work contributes positively to developing trustworthy and interpretable language models.
Acknowledgments
This work was partially supported by the National Research Foundation of Korea (NRF) under Grant No. RS-2023-00222663 (Center for Optimizing Hyperscale AI Models and Platforms), and by the Institute for Information and Communications Technology Promotion (IITP) under Grant No. 2018-0-00581 (CUDA Programming Environment for FPGA Clusters) and No. RS-2025-02304554 (Efficient and Scalable Framework for AI Heterogeneous Cluster Systems), all funded by the Ministry of Science and ICT (MSIT) of Korea. Additional support was provided by the BK21 Plus Program for Innovative Data Science Talent Education (Department of Data Science, SNU, No. 5199990914569), funded by the Ministry of Education (MOE) of Korea. This work was also partially supported by Samsung Display Co., Ltd. and the Artificial Intelligence Industrial Convergence Cluster Development Project, funded by the MSIT and Gwangju Metropolitan City. Research facilities were provided by ICT at Seoul National University.
Appendix A Copular Verbs
Copular verbs, also known as linking verbs, are verbs that connect the subject of a sentence to a subject complement, which can be a noun, adjective, or other expression that describes or identifies the subject. Unlike action verbs, copular verbs do not express actions but rather states or conditions. The most common copular verb in English is "to be" in its various forms (am, is, are, was, were). Other examples include "seem," "appear," "become," "feel," "look," "sound," "taste," and "smell" when used to describe the subject’s state (Hengeveld, 1986).
In the context of standard negation, as discussed in Section 3.3, negation typically targets the main predicate of a clause. For sentences with copular verbs, this means that the entire verb phrase, including the copular verb and its complement, is subject to negation. For example, in the sentence "She is a doctor," the main predicate is "is a doctor." Negating this sentence results in "She is not a doctor," where the negation applies to the entire predicate, not just the verb "is."
This approach aligns with the semantic characterization of standard negation, where the truth value of the main predicate is reversed ($P$rightarrow $\neg P$). In copular constructions, the negation encompasses both the copular verb and its complement, effectively altering the state or identity attributed to the subject.
In the case of copular verbs, standard negation targets the entire predicate, including the complement. This means that negation can be realized either syntactically (e.g., "is not an expert") or by replacing the complement with its complementary antonym (e.g., "is a non-expert"), both of which result in the reversal of the main predicate’s truth value. Although such constructions may superficially appear to be non-verbal negation, especially when the complement is a noun or adjective, they are, in fact, instances of verbal negation, since the negation applies to the predicate as a whole.

### v3

@@ Appendix A Typology of Contradiction
Table 7: Contradiction types from de2008finding. Contradiction covers a broader scope than negation.
Appendix B Copular Verbs
Copular verbs, also known as linking verbs, are verbs that connect the subject of a sentence to a subject complement, which can be a noun, adjective, or other expression that describes or identifies the subject. Unlike action verbs, copular verbs do not express actions but rather states or conditions. The most common copular verb in English is "to be" in its various forms (am, is, are, was, were). Other examples include "seem," "appear," "become," "feel," "look," "sound," "taste," and "smell" when used to describe the subject’s state (hengeveld1986copular).
As discussed in Section 3.3, standard negation in this work targets the main predicate of a clause. For sentences with copular verbs, this means that the entire verb phrase, including the copular verb and its complement, is subject to negation. For example, in the sentence "She is a doctor," the main predicate is "is a doctor." Negating this sentence results in "She is not a doctor," where the negation applies to the entire predicate, not just the verb "is."
Negation of a verb phrase including a copular verb can be realized either syntactically (e.g., "is not an expert") or by replacing the complement with its complementary antonym (e.g., "is a non-expert"), both of which result in the reversal of the main predicate’s truth value. Although such constructions may superficially appear to be non-verbal negation, especially when the complement is a noun or adjective, they are, in fact, instances of verbal negation, since the negation applies to the predicate as a whole.
Appendix C Negation of Implications
Negating implications presents challenges, as natural language intuitions often diverge from the rules of formal logic. Let’s say there is a conditional statement, "If I study hard, I will pass the bar exam." Formally, let $P$ denote "I study hard" and $Q$ denote "I will pass the bar exam." In classical logic, the conditional "if $P$, QQ" can be false only when $P$ is true and $Q$ is false. This implies that the negation of the conditional is $P\;\text{and}\;\text{Neg}(Q)$" ("I study hard and I won’t pass the bar exam,) while the conditional itself is equivalent to $\text{Neg}(P)\;\text{or}\;Q$" ("I don’t study hard or I will pass the exam) (nguyen2023negation).
Psychological studies confirm that people often accept both "if $P$, Neg(QQ)" ("If I study hard, I won’t pass the exam.") and "if Neg($P$), QQ" ("If I don’t study hard, I will pass the exam."). However, the former can be interpreted as $\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$", and the latter $P\;\text{or}\;Q$", both of which are not equivalent to the original statement’s negation, $P\;\text{and}\;\text{Neg}(Q)$". "if Neg($P$), Neg(QQ)" ("If I don’t study hard, I won’t pass the exam.") is not the correct negation as well, as it is equivalent to $P\;\text{or}\;\text{Neg}(Q)$(espino2012not).
While humans often struggle to distinguish the correct negation of a conditional from invalid alternatives, the logical form is unambiguous. We therefore include conditional statements in our benchmark to test whether language models, like humans, are prone to intuitive but invalid interpretations, or whether they can correctly apply truth-functional reasoning.
Note that implications in natural language are not limited to the explicit "if $P$, QQ" form, but may also appear with connectives such as when, as long as, or unless, which functionally convey conditional meaning and are treated under the same negation principle.
Appendix D Compound Sentences and Coordinating Conjunction
A compound sentence consists of two or more independent clauses joined by a coordinating conjunction. Each clause can stand alone, but they are combined to express related ideas (gleitman1965coordinating).
  • "She studied hard, and she passed the exam."
  • "I wanted to go, but it was raining."
  • "You can call me, or you can send an email."
We consider only the coordinating conjunctions and, or, and but as indicators of compound sentences, in which two or more independent clauses are equally connected. Although but introduces a contrast semantically, in terms of logical structure, it functions as a conjunction equivalent to and; therefore, its negation follows the same principle.
Appendix E Local Negation Constructions Excluded
In constructing the Thunder-NUBench dataset, we consider various types of local (i.e., subclausal) negation, where negation applies to a phrase or constituent rather than the main predicate. However, several constructions are excluded due to their semantic ambiguity, syntactic irregularity, or misalignment with the benchmark’s focus on verbal negation.
Infinitive Phrase Negation.
Infinitive phrases (e.g., "to go") can be negated with "not" (e.g., "not to go" or "to not go"). Unlike the clause-level structures that define our local negation category, infinitive phrases are not full clauses but simply part of a verb phrase, making them less compatible with our definition. Moreover, although grammatically correct, this construction is relatively rare and sounds awkward depending on the context.
  • Original: George wants to go to the park.
  • Negated (infinitive): George wants not to go / George wants to not go to the park.
For these reasons, we exclude infinitive phrase negation from the benchmark.
Appositive Clause Negation.
  • Original: My brother, a talented musician, plays the guitar.
Such changes alter descriptive content rather than reversing the meaning of the predicate, and often fall into the domain of contradiction. Accordingly, they are excluded from the dataset.
Prepositional Phrase Negation.
  • Original: She went to the park with her bird.
  • Negated (preposition): She went to the park without her bird.
In all of the above cases, the negation does not target the whole verb phrase but rather peripheral elements within the sentence. As the Thunder-NUBench is designed to evaluate verbal negation, these local or phrase-level forms of negation were intentionally left out.
Appendix F Double Negation
Double negation refers to the use of two forms of grammatical negation within a single sentence. In standard English, only one negative form should be present in a subject-predicate construction; the presence of two negatives is generally considered non-standard and often results in an unintended meaning. For example, while "He’s going nowhere" is correct, "He’s not going nowhere" is ungrammatical. Another example is "I won’t bake no cake," which combines verb negation ("won’t") with object negation ("no cake"), resulting in a grammatically incorrect construction (deprez2015double).
In English, certain double negation constructions convey affirmative meanings rather than intensifying negation, effectively paraphrasing the original positive statement (e.g., $\neg\neg p\approx p$(van1996litotes). This rhetorical device, known as litotes, often manifests in expressions such as "not bad," implying "good," or "not unhappy," implying "happy." Leveraging this phenomenon, we have generated paraphrase candidates for our dataset using such double negation patterns. For example,
However, upon closer examination, these paraphrase candidates do not always preserve the exact meaning of the original sentence. The antonyms used (e.g., "exclude" for "cover," "unhappy" for "honoured") are not always true complementary antonyms, which does not effectively negate the meaning. Moreover, the litotes construction ("does not fail to fuse") tends to add an emphatic nuance, rather than being a perfect semantic equivalent. Therefore, the boundary between paraphrasing and double negation is ambiguous, and their relationship requires more careful analysis. Given these issues, and because our primary focus is on standard negation, we ultimately decide to exclude double negation constructions as paraphrase candidates from our dataset.
Appendix G HoVer Dataset